Walnut Systems produces two different products, Product A, which sells for $220 per unit, and Product B, which sells for $280 per unit, using three different activities: Design, which uses Engineering Hours as an activity driver; Machining, which uses machine hours as an activity driver; and Inspection, which uses number of batches as an activity driver. The cost of each activity and usage of the activity drivers are as follows:

Cost Usage by Product A Usage by Product B
Design (Engineering hours) $ 150,000 300 200
Machining (Machine Hours) $ 500,000 1,400 2,600
Inspection (Batches) $ 30,000 15 45


Walnut manufactures 10,000 units of Product A and 8,150 units of Product B per month. Each unit of Product A uses $30 of direct materials and $40 of direct labor, while each unit of Product B uses $85 of direct materials and $45 of direct labor.

Required:
a.
Calculate the Design activity proportions for Products A and B.




b. Calculate the Machining activity proportions for Products A and B.



c. Calculate the Inspection activity proportions for Products A and B.



d. Determine the indirect costs assigned to Product A.



e. Determine the indirect costs assigned to Product B.



f. Determine the manufacturing cost per unit for Product A. (Round your answer to 2 decimal places.)



g. Determine the manufacturing cost per unit for Product B.



h. Determine the gross profit per unit for Product A. (Round your intermediate calculation to 2 decimal places. Round your answer to 2 decimal places.)


i. Determine the gross profit per unit for Product B.
Answer #1

a. Design Activity Proportions
Product A = 300/500 = 60%
Product B = 200/500 = 40%

b. Machining Activity Proportions
Product A = 1400/4000 = 35%
Product B = 2600/4000 = 65%

c. Inspection Activity Proportions
Product A = 15/60 = 25%
Product B = 45/60 = 75%

d. Product A

Indirect Costs
Design $        90,000
Machining $     1,75,000
Inspection $           7,500
Total Indirect Costs $     2,72,500
Units 10000
Indirect Cost per unit $          27.25

e. Product B

Indirect Costs
Design $        60,000
Machining $     3,25,000
Inspection $        22,500
Total Indirect Costs $     4,07,500
Units 8150
Indirect Cost per unit $          50.00

f-i

Product A Product B
Direct Material $          30.00 $          85.00
Direct Labor $          40.00 $          45.00
Indirect Cost $          27.25 $          50.00
Manufacturing Cost per unit $          97.25 $        180.00
Selling Price $        220.00 $        280.00
Gross Profit per unit $        122.75 $        100.00